Solve for ‘x’ and ‘y’ given the following two equations using the elimination method:

2x + 3y = 8
x = 2y - 3

Answer by TimothyLamb(4379)

2x + 3y = 8
x = 2y - 3
---
2x + 3y = 8
x - 2y = -3
---
2x + 3y = 8
-2x + 4y = 6
---
add:
7y = 14
---
y = 2
x = 1
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php